OpenStudy (anonymous):

What is a word or story problem that do solves the equation 4y-7=3y

OpenStudy (anonymous):

are you gonna give us choices or do you want us to give you a word problem?

OpenStudy (anonymous):

collegegirl92 has 3 times fs collegegirl92 friend has 4 times fs - 7 how many fs do they have? Something like that? I could be wrong tho

OpenStudy (anonymous):

Idk lol

OpenStudy (anonymous):

It's a simple speed problem. Or purchasing some thing that costs you a montly bill

OpenStudy (anonymous):

I neef a word problem that can be solve by this equation

OpenStudy (anonymous):

like you can say. You wanna get product A. Store B rents it for $3 a month. Or store C has it for $4 a month put gives you $7 cash back howmany months until both cost you the same? the equation will be 4y-7=3y

OpenStudy (anonymous):

Ok I need help figuring out how to do it on my own

OpenStudy (anonymous):

those are linear problems. 4y-7=3y are of type y=mx+b you get int in a y=mx+b form by solving for y. THose occur alot in finance and physics. In physics they are velocity problems There are also sort of age problems where they say amy is x times older than sam and sam is y times plus 7 years older. THey all involve the same y=mx+b formula to model it.